OpenStudy (ageag):

There are (72)3 ⋅ 70 lambs on a farm. What is the total number of lambs on the farm? 0 7 75 76

OpenStudy (ageag):

The 2, 3 and 0 are small

OpenStudy (ageag):

@wckeller

OpenStudy (wckeller):

like this? \[(7^2)^3*7^0\]

OpenStudy (ageag):

Yeah

OpenStudy (wckeller):

\[7^2 = 49\] \[49^3=117,649\] \[7^0=1\] \[117,649*1=117,649\] Final answer= 117,649